OpenStudy (anonymous):

what will be the change in value of xy/(x^3) when x is reduced by 50%

OpenStudy (anonymous):

@amistre64 any idea on how to go about this one?

ganeshie8 (ganeshie8):

\(xy/x^3 = y/x^2\)

ganeshie8 (ganeshie8):

call that f(x) f(x) = y/x^2 -----(1)

ganeshie8 (ganeshie8):

f(.5x) = y/(.5x)^2 = y/.25(x^2)------- (2)

ganeshie8 (ganeshie8):

take (1)/(2) to find the change in expression

ganeshie8 (ganeshie8):

f(x) / f(.5x) = ?

OpenStudy (anonymous):

@ganeshie8 0.25=25%

ganeshie8 (ganeshie8):

yeah it reduces the value of expression to 25 % or f(x) reduces by 75 % both are same

OpenStudy (anonymous):

Thank you

ganeshie8 (ganeshie8):

np